# Optimized Memory Tagging on AmpereOne® Processors

### v2

@@ V-A Tag Storage Considerations
In server platforms populated with memory technology that has ECC support, one option to store allocation tags is to use a subset of the ECC bits. Since these bits are used only by the memory controller unit (MCU) to provide memory reliability features, use of these bits does not impact total memory capacity available to software and eliminates the impact on VM density and TCO degradation to the cloud service provider. AmpereOne® SoC servers use DDR5 registered DIMMs with ECC support providing access of 80 Byte Codewords (64B data + 16B of parity ECC), with a capability to correct a full symbol of 1 Byte. While this allows use of the ECC bits for tag storage, micro-architecture enhancements are needed in the MCU to provide the memory reliability capabilities needed in datacenter platforms using the remainder of the ECC bits available. This support is discussed later in section VI-B. A similar technique for storing metadata in ECC bits is found in architectures for Confidential Computing [cryptoeprint:2022/1472] and GPU memory safety [sullivan2023implicit].
@@ V-B Tag Fetching Considerations
The organization of MTE tags in memory also affects the tag fetch overhead and hence, the run-time performance of applications that use MTE. If tags are not co-located with their data, the CPU read requests to tagged memory will require separate reads to fetch the corresponding allocation tags; these tag reads consume additional memory bandwidth and also potentially increase memory latency for read operations. CPU reads to untagged memory will not require additional memory reads but may be indirectly affected by the increased traffic due to tagged memory reads. Co-locating tag storage with data allows reading the tags along with the data without generating additional memory traffic and hence, doesn’t negatively impact memory bandwidth and latency. Figure 2 shows the relative impact on bandwidth consumption of the MCU and Mesh interconnect on each core cache miss, based on the choice of tag storage method. RMW indicates that the memory subsystem is required to perform a Read-Modify-Write to update the MTE tags. Previous work has demonstrated the performance impact of the increased traffic for memory tagging, in both CPU and GPU architectures [lamster2024voodoo, sullivan2023implicit, cryptoeprint:2022/1472].
@@ VI-B MTE in the MCU and DRAM
There is no latency impact within the MCU, as the tag bits are transferred along with the data bits, similar to Error-Correcting Code (ECC) bits. However, a notable impact arises concerning the utilization of ECC for error detection and correction. In certain DRAM and ECC configurations, insufficient metadata bits may be available to store MTE tags. To address this, some ECC bits must be re-purposed for tag storage. Ampere accommodates this by offering multiple ECC schemes tailored to various DRAM configurations. Specifically, a novel ECC scheme is employed that provides the requisite reliability protection for datacenter class SoCs while simultaneously making bits available for MTE allocation tags [ecc2022]. The selection of the appropriate ECC scheme is determined by both the DIMM configuration and the desired MTE support. AmpereOne® SoC systems support four schemes for ECC:
@@ VII-A Platform Configuration
The System Under Test (SUT) utilized an Ampere Computing reference platform, designated as Mt. Mitchell [ampere_mtmitchell]. This rack server was equipped with an AmpereOne® M SoC, featuring 192 cores operating at 3.2 GHz. Memory consisted of 512 GB of SK Hynix DDR5 modules, arranged in an 8x64 GB configuration and operating at 5200 MT/s. For I/O, Samsung NVMe Solid State Drives (SSDs) provided storage, and Mellanox ConnectX 100GbE cards managed all network communications. To generate traffic, two 128-core Ampere Altra® Max servers served as client systems. The clients are configured to keep the servers fully active under a P99 service level agreement. Typically, there is a one-to-one correspondence between client and SUT cores to saturate the system for performance measurement.
@@ VII-B Software Environment and Workloads
The Linux kernel has supported MTE since version 5.1.0 [Linux_MTE], with corresponding support added to GNU Binutils in version 2.45 [GNU_Binutils] and to the GNU C Library (glibc) in version 2.33 [Glib_MTE]. Fedora 36 and later distributions ship their kernels with CONFIG_ARM64_MTE enabled by default. For the following experiments, the system under test operated on Fedora 40 with GNU/Linux 6.10.6 and glibc 2.39.
To assess MTE’s performance impact, motivated by similar evaluations [utaustin_mte], data was collected from a diverse range of real-world datacenter applications. These are industry-standard benchmarks, widely recognized for datacenter workload characterization and competitive performance evaluation. In all benchmark runs, all available cores on the SoC were fully utilized, to the extent each benchmark allowed. A summary of the benchmarks utilized is provided below.
memcached / memtier [memcached, memtier]: memcached is an open source in-memory NoSQL key/value distributed object caching database application. For performance evaluation, a client-server configuration was employed, using memtier_benchmark-v1.3.0 load generators connected over a high-speed network to the server SUT. The SUT executed multiple independent instances of memcached-v1.6.21.
@@
Redis / memtier [redis, memtier]: Redis is an open source in-memory NoSQL key/value data store that functions as either a database or an application cache. The experimental setup employed a client-server configuration: multiple client systems, running the memtier_benchmark-v1.3.0 load generation tool, connected over a high-speed network to the System Under Test. The SUT executes multiple, independent, single-threaded instances of Redis-v7.2.0 servers. While Redis workloads typically leverage the jemalloc memory allocation library [jemalloc_paper], jemalloc currently does not support memory tagging. Therefore, to evaluate tagging impact, Redis was linked with glibc to use the standard malloc memory allocator.
vbench (H.264 Video Transcoding Benchmark) [vbench]: vbench is a benchmark designed for evaluating H.264 video transcoding scenarios in cloud-based video-as-a-service (VaaS) applications. It assesses transcoding performance relevant to typical cloud video workflows, and employs 15 H.264-compressed input files, varying in resolution (480p to 4K) and frame rates (25 to 60 fps). For this analysis, the focus was on two distinct VaaS profiles, Upload and Video On Demand (VOD):
@@
nginx / wrk [nginx, wrkgit]: nginx is an open-source web server commonly employed as a reverse proxy, load balancer, or HTTP cache. To focus on server-side computational overhead, the SUT was configured with Brotli compression and the workload requests initiated server-side Lua scripts [cloudflare]. Given this CPU-bound scenario, a one-to-one client-server configuration is employed where a single client running the wrk-v4.1.0 load generator was sufficient to saturate the SUT. The client connected via a high-speed network to the SUT, which ran a single instance of nginx-v1.24 that spawned a worker process per core.
MySQL / sysbench [mysql, sysbench]: MySQL is an open source, SQL-compliant, relational database management system (RDBMS). For performance evaluation Sysbench v1.1, a multi-threaded load generation and benchmarking tool, is employed. Sysbench was used to establish a simple database schema, populate database tables, and generate multi-threaded SQL query workloads. These queries were directed to a single instance of the MySQL-v8.0.43 database server, running concurrently with Sysbench on the SUT and communicating via TCP over a local network interface.
PostgreSQL / HammerDB [postgresql, hammerdb]: PostgreSQL is an open source SQL-compliant, object-relational database management system (ORDBMS). For performance evaluation a client system running HammerDB-v4.3, a multi-threaded load generation and benchmarking tool to generate a TPC-C-like workload, was employed. This workload was transmitted over a high-speed network to the SUT, executing a single instance of the PostgreSQL v15.3 database server.
SPEC CPU® 2017 integer base  [cpu2017]: SPEC CPU was run in refrate mode with 192 copies to saturate the system and estimate SPECrate®2017_int_base. The binaries were built with open-source community gcc-15.1.1 with these base optimization flags: -O3 -mcpu=ampere1a -flto.
@@ VII-C Performance Analysis
Contemporary research [utaustin_mte] has analyzed ARM MTE performance overheads using single-copy SPEC CPU, and a consolidated multi-tenant setup where both the server benchmark and client load generators are hosted on a single machine. In contrast, the measurements presented below are from 192-copy SPEC CPU and fully utilized server platforms with discrete client load generators. While both approaches are valid, this one closely emulates real-world cloud deployments operating at peak utilization, and offers a unique yet complementary perspective on MTE performance.
@@
To guide these two decisions, both performance ratios across the benchmarks are shown. Figure 4 presents the performance ratios and measurement metrics for each datacenter cloud workload. Figure 5 reports the ratios for each SPEC CPU 2017 benchmark in 192-copy refrate. To account for the inherent run-to-run variability observed in multi-threaded cloud workloads, each benchmark was executed five times. The resulting data were statistically analyzed using ministat [ministat], which compares two small populations to compute performance ratios and their associated error bars using Student’s t-test [studentt]. All statistical comparisons were performed at a 99% confidence level. Results designated as "on par" indicate that no statistically significant performance difference could be discerned between the compared populations.
When looking at the B vs. A comparisons, enabling MTE in hardware can introduce a small overhead even when applications are not using tagging, due to SoC mechanisms that preserve tag state through the memory hierarchy. This platform-level comparison is relevant for multi-tenant deployments where a cloud provider enables MTE across the fleet. Across the workloads evaluated in Figures 4 and 5, that overhead was in the 1–6% range, while memcached exhibited a 7% improvement (with high variance). In this case, tag-loads effectively become prefetches, since tag-checking is turned off. Since tags and data are co-located, cache lines are brought in to the L1 data cache which happen to be beneficial to the instruction stream being executed. Practically, these numbers suggest that tenants not asking for tag checking experience only modest impact when the platform enables MTE support for other users.
The primary sources of hardware slowdown stem from first-generation MTE integration in the AmpereOne® microarchitecture and pipelines. First, as also observed by [utaustin_mte], store-to-load forwarding opportunities are reduced for tag-checked stores in the core. Second, the additional tag-read traffic increases pressure on L1D cache read ports, introducing structural hazards in the load/store pipeline. Refinements to address both of these issues have been incorporated into the next generation of Ampere’s CPU cores.
@@
Second, at the kernel interface, glibc disables use of brk(2) for its arenas when tagging is requested and instead relies on mmap(2). The brk(2) path, which primarily advances the program break to grow the heap (a simple move of a top-of-heap pointer), does not readily provide page attributes or initialization required for MTE such as setting PROT_MTE and preparing tag storage. In contrast, mmap(2) allows the allocator to request appropriately protected anonymous mappings, but it incurs higher instruction overhead: selecting a suitable virtual-address range, creating or merging VMAs (Virtual Memory Areas), updating page tables, honoring protection and flags, and triggering first-touch work and tag initialization. The kernel marks these pages as MTE-capable, and tag setup further increases instruction count and latency on first access.
@@
This hypothesis aligns with the observed behavior. Workloads like 502.gcc, which construct complex, graph-like data structures (e.g., abstract syntax trees), naturally produce a more disjoint and fragmented allocation pattern. This fragmentation is exacerbated by 502.gcc’s frequent use of realloc(3). When these reallocations require moving data, their cost is amplified under MTE: the old memory region must be de-tagged and the new region must be tagged, effectively doubling the tag management overhead for a single high-level operation. In contrast, the more linear, stream-like processing in 500.perlbench with zero realloc’s likely results in better memory locality and a less fragmented VMA layout, mitigating some of the kernel- and hardware-level penalties despite also having a high allocation rate of small objects. This finding is consistent with studies of other hardware-based memory safety features; prior research [kim_heap_safety] similarly identified the high frequency of heap operations in 502.gcc and 520.omnetpp as the primary performance limiter due to the overhead of extra operations added to the allocator’s critical path.
@@
To further validate the hypothesis that frequent, fragmented allocations are a primary source of overhead, we evaluated the performance of these workloads with jemalloc [jemalloc_paper], a memory allocator designed to mitigate fragmentation. It accomplishes this through techniques such as size-classed arenas, which are particularly effective for workloads with high rates of small-object allocations. If our hypothesis is correct, the applications impacted by MTE should, in turn, derive benefit from a fragmentation-aware allocator like jemalloc.
Our results confirm this correlation. In Figure 8, we show jemalloc’s performance improvement on 192-copy CPU 2017, with a tagging-disabled baseline (since jemalloc does not support tagging). The workloads that exhibited the largest MTE-related regressions (502.gcc, 523.xalanc, and 520.omnetpp) do gain substantial performance when linked and run with jemalloc. This finding provides strong evidence that the observed MTE slowdowns are not an intrinsic cost of tag checking itself, but are tightly coupled to the underlying memory allocation patterns and the efficiency of the memory management subsystem. This study provides motivation for the community to research custom heap allocators with tagging support, which can additionally assist with other types of security issues [alt_alloc, scudo]. Using type-based allocators or size-based allocators also aligns with Apple’s direction of using a custom allocator for MTE [AppleMemoryIntegrityEnforcementBlog].

### v3

@@ V-A Tag Storage Considerations
In server platforms populated with memory technology that has ECC support, one option to store allocation tags is to use a subset of the ECC bits. Since these bits are used only by the memory controller unit (MCU) to provide memory reliability features, use of these bits does not impact total memory capacity available to software and eliminates the impact on VM density and TCO degradation to the cloud service provider. AmpereOne® SoC servers use DDR5 registered DIMMs with ECC support providing access of 80 Byte Codewords (64B data + 16B of parity ECC), with a capability to correct a full symbol of 1 Byte. While this allows use of the ECC bits for tag storage, micro-architecture enhancements are needed in the MCU to provide the memory reliability capabilities needed in datacenter platforms using the remainder of the ECC bits available. This support is discussed later in Section VI-B. A similar technique for storing metadata in ECC bits is found in the Sun SPARC ADI implementation for Memory Tagging [67], the RISC-V based Rocket SOC from lowRISC.org [66], architectures for Confidential Computing [13] and GPU memory safety [68].
@@ V-B Tag Fetching Considerations
The organization of MTE tags in memory also affects the tag fetch overhead and hence, the run-time performance of applications that use MTE. If tags are not co-located with their data, the CPU read requests to tagged memory will require separate reads to fetch the corresponding allocation tags; these tag reads consume additional memory bandwidth and also potentially increase memory latency for read operations. CPU reads to untagged memory will not require additional memory reads but may be indirectly affected by the increased traffic due to tagged memory reads. Co-locating tag storage with data allows reading the tags along with the data without generating additional memory traffic and hence, doesn’t negatively impact memory bandwidth and latency. Figure 4 shows the relative impact on bandwidth consumption of the MCU and Mesh interconnect on each core cache miss, based on the choice of tag storage method. RMW indicates that the memory subsystem is required to perform a Read-Modify-Write to update the MTE tags. Previous work has demonstrated the performance impact of the increased traffic for memory tagging, in both CPU and GPU architectures [42, 68, 13].
@@ VI-B MTE in the MCU and DRAM
There is no latency impact within the MCU, as the tag bits are transferred along with the data bits, similar to Error-Correcting Code (ECC) bits. However, a notable impact arises concerning the utilization of ECC for error detection and correction. In certain DRAM and ECC configurations, insufficient metadata bits may be available to store MTE tags. To address this, some ECC bits must be re-purposed for tag storage. Ampere accommodates this by offering multiple ECC schemes tailored to various DRAM configurations. Specifically, a novel ECC scheme is employed that provides the requisite reliability protection for datacenter class SoCs while simultaneously making bits available for MTE allocation tags [69]. The selection of the appropriate ECC scheme is determined by both the DIMM configuration and the desired MTE support. AmpereOne® SoC systems support four schemes for ECC:
@@ VII-A Platform Configuration
The System Under Test (SUT) utilized an Ampere Computing reference platform, designated as Mt. Mitchell [4]. This rack server was equipped with an AmpereOne® M SoC, featuring 192 cores operating at 3.2 GHz. Memory consisted of 512 GB of SK Hynix DDR5 modules, arranged in an 8x64 GB configuration and operating at 5200 MT/s. The configuration supported SymbolECC for RAS capabilities. For I/O, Samsung NVMe Solid State Drives (SSDs) provided storage, and Mellanox ConnectX 100GbE cards managed all network communications. To generate traffic, two 128-core Ampere Altra® Max servers served as client systems. The clients are configured to keep the servers fully active under a P99 service level agreement. Typically, there is a one-to-one correspondence between client and SUT cores to saturate the system for performance measurement.
@@ VII-B Software Environment and Workloads
The Linux kernel has supported MTE since version 5.1.0 [73], with corresponding support added to GNU Binutils in version 2.45 [50] and to the GNU C Library (glibc) in version 2.33 [49]. Fedora 36 and later distributions ship their kernels with CONFIG_ARM64_MTE enabled by default. For the following experiments, the system under test operated on Fedora 40 with GNU/Linux 6.10.6 and glibc 2.39.
To assess MTE’s performance impact, motivated by similar evaluations [55], data was collected from a diverse range of real-world datacenter applications. These are industry-standard benchmarks, widely recognized for datacenter workload characterization and competitive performance evaluation. In all benchmark runs, all available cores on the SoC were fully utilized, to the extent each benchmark allowed.
Memory tagging-specific overheads for tag fetching and checking are identical for workloads running bare-metal or within a VM; no additional VM-exits are introduced by tag setting or checking. We measured the MTE performance in a VM for subset of the workloads early in the AmpereOne program and found that the virtualization overhead (primarily from stage-2 translations) slightly dilutes the MTE performance impact compared to a bare-metal configuration. Hence, the paper focuses on the worse case of the two, bare-metal workloads, to fully expose the performance impact.
A summary of the benchmarks utilized is provided below.
memcached / memtier [48, 56]: memcached is an open source in-memory NoSQL key/value distributed object caching database application. For performance evaluation, a client-server configuration was employed, using memtier_benchmark-v1.3.0 load generators connected over a high-speed network to the server SUT. The SUT executed multiple independent instances of memcached-v1.6.21.
@@
Redis / memtier [61, 56]: Redis is an open source in-memory NoSQL key/value data store that functions as either a database or an application cache. The experimental setup employed a client-server configuration: multiple client systems, running the memtier_benchmark-v1.3.0 load generation tool, connected over a high-speed network to the System Under Test. The SUT executes multiple, independent, single-threaded instances of Redis-v7.2.0 servers. While Redis workloads typically leverage the jemalloc memory allocation library [26], jemalloc currently does not support memory tagging. Therefore, to evaluate tagging impact, Redis was linked with glibc to use the standard malloc memory allocator.
vbench (H.264 Video Transcoding Benchmark) [47]: vbench is a benchmark designed for evaluating H.264 video transcoding scenarios in cloud-based video-as-a-service (VaaS) applications. It assesses transcoding performance relevant to typical cloud video workflows, and employs 15 H.264-compressed input files, varying in resolution (480p to 4K) and frame rates (25 to 60 fps). For this analysis, the focus was on two distinct VaaS profiles, Upload and Video On Demand (VOD):
@@
nginx / wrk [54, 27]: nginx is an open-source web server commonly employed as a reverse proxy, load balancer, or HTTP cache. To focus on server-side computational overhead, the SUT was configured with Brotli compression and the workload requests initiated server-side Lua scripts [41]. Given this CPU-bound scenario, a one-to-one client-server configuration is employed where a single client running the wrk-v4.1.0 load generator was sufficient to saturate the SUT. The client connected via a high-speed network to the SUT, which ran a single instance of nginx-v1.24 that spawned a worker process per core.
MySQL / sysbench [53, 40]: MySQL is an open source, SQL-compliant, relational database management system (RDBMS). For performance evaluation Sysbench v1.1, a multi-threaded load generation and benchmarking tool, is employed. Sysbench was used to establish a simple database schema, populate database tables, and generate multi-threaded SQL query workloads. These queries were directed to a single instance of the MySQL-v8.0.43 database server, running concurrently with Sysbench on the SUT and communicating via TCP over a local network interface.
PostgreSQL / HammerDB [58, 33]: PostgreSQL is an open source SQL-compliant, object-relational database management system (ORDBMS). For performance evaluation a client system running HammerDB-v4.3, a multi-threaded load generation and benchmarking tool to generate a TPC-C-like workload, was employed. This workload was transmitted over a high-speed network to the SUT, executing a single instance of the PostgreSQL v15.3 database server.
SPEC CPU® 2017 integer base  [16]: SPEC CPU was run in refrate mode with 192 copies to saturate the system and estimate SPECrate®2017_int_base. The binaries were built with open-source community gcc-15.1.1 with these base optimization flags: -O3 -mcpu=ampere1a -flto.
@@ VII-C Performance Analysis
Contemporary research [55] has analyzed ARM MTE performance overheads using single-copy SPEC CPU, and a consolidated multi-tenant setup where both the server benchmark and client load generators are hosted on a single machine. In contrast, the measurements presented below are from 192-copy SPEC CPU and fully utilized server platforms with discrete client load generators. While both approaches are valid, this one closely emulates real-world cloud deployments operating at peak utilization, and offers a unique yet complementary perspective on MTE performance.
@@
To guide these two decisions, both performance ratios across the benchmarks are shown. Figure 6 presents the performance ratios and measurement metrics for each datacenter cloud workload. Figure 7 reports the ratios for each SPEC CPU 2017 benchmark in 192-copy refrate. To account for the inherent run-to-run variability observed in multi-threaded cloud workloads, each benchmark was executed five times. The resulting data were statistically analyzed using ministat [36], which compares two small populations to compute performance ratios and their associated error bars using Student’s t-test [31]. All statistical comparisons were performed at a 99% confidence level. Results designated as “on par” indicate that no statistically significant performance difference could be discerned between the compared populations.
@@
When looking at the B vs. A comparisons, enabling MTE in hardware can introduce a small overhead even when applications are not using tagging, due to SoC mechanisms that preserve tag state through the memory hierarchy. This platform-level comparison is relevant for multi-tenant deployments where a cloud provider enables MTE across the fleet. Across the workloads evaluated in Figures 6 and 7, that overhead was in the 1–6% range, while memcached exhibited a 7% improvement (with high variance). In this case, tag-loads effectively become prefetches, since tag-checking is turned off. Since tags and data are co-located, cache lines are brought in to the L1 data cache which happen to be beneficial to the instruction stream being executed. The source of the MTE performance benefit to memcached was confirmed with PMU data, which showed a significant reduction in time spent waiting for L2 misses, without a significant change in other operations which can prefetch lines into the cache, such as hardware prefetches or wrong-path speculation. Practically, these numbers suggest that tenants not asking for tag checking experience only modest impact when the platform enables MTE support for other users.
@@
The primary sources of hardware slowdown stem from first-generation MTE integration in the AmpereOne® microarchitecture and pipelines. First, as also observed by [55], store-to-load forwarding opportunities are reduced for tag-checked stores in the core. Second, the additional tag-read traffic increases pressure on L1D cache read ports, introducing structural hazards in the load/store pipeline. Refinements to address both of these issues have been incorporated into the next generation of Ampere’s CPU cores.
@@
Second, at the kernel interface, glibc disables use of brk(2) for its arenas when tagging is requested and instead relies on mmap(2). The brk(2) path, which primarily advances the program break to grow the heap (a simple move of a top-of-heap pointer), does not readily provide the page attributes or initialization required for MTE such as setting PROT_MTE and preparing tag storage. In contrast, mmap(2) allows the allocator to request appropriately protected anonymous mappings, but it incurs higher instruction overhead: selecting a suitable virtual-address range, creating or merging VMAs (Virtual Memory Areas), updating page tables, honoring protection and flags, and triggering first-touch work and tag initialization. The kernel marks these pages as MTE-capable, and tag setup further increases instruction count and latency on first access.
@@
This hypothesis aligns with the observed behavior. Workloads like 502.gcc, which construct complex, graph-like data structures (e.g., abstract syntax trees), naturally produce a more disjoint and fragmented allocation pattern. This fragmentation is exacerbated by 502.gcc’s frequent use of realloc(3). When these reallocations require moving data, their cost is amplified under MTE: the old memory region must be de-tagged and the new region must be tagged, effectively doubling the tag management overhead for a single high-level operation. In contrast, the more linear, stream-like processing in 500.perlbench with zero realloc’s results in better memory locality and a less fragmented VMA layout, mitigating some of the kernel- and hardware-level penalties despite also having a high allocation rate of small objects. This finding is consistent with studies of other hardware-based memory safety features; prior research [38] similarly identified the high frequency of heap operations in 502.gcc and 520.omnetpp as the primary performance limiter due to the overhead of extra operations added to the allocator’s critical path.
@@
To further validate the hypothesis that frequent, fragmented allocations are a primary source of overhead, we evaluated the performance of these workloads with jemalloc [26], a memory allocator designed to mitigate fragmentation. It accomplishes this through techniques such as size-classed arenas, which are particularly effective for workloads with high rates of small-object allocations. If our hypothesis is correct, the applications impacted by MTE should, in turn, derive benefit from a fragmentation-aware allocator like jemalloc.
Our results confirm this correlation. In Figure 10, we show jemalloc’s performance improvement on 192-copy CPU 2017, with a tagging-disabled baseline (since jemalloc does not support tagging). The workloads that exhibited the largest MTE-related regressions (502.gcc, 523.xalanc, and 520.omnetpp) do gain substantial performance when linked and run with jemalloc. This finding provides strong evidence that the observed MTE slowdowns are not an intrinsic cost of tag checking itself, but are tightly coupled to the underlying memory allocation patterns and the efficiency of the memory management subsystem. This study provides motivation for the community to research custom heap allocators with tagging support and smarter policies, which can additionally assist with other types of security issues [24, 45]. Using type-based allocators or size-based allocators also aligns with Apple’s direction of using a custom allocator for MTE, xzone malloc [7], where small blocks’ tags are reassigned on free but larger blocks’ tags are reassigned lazily upon the next allocation [65]. The GNU Tools Cauldron 2025 sessions also acknowledged the issue with MTE and GLIBC malloc of small objects [19].